# Exploring Applications of State Space Models and Advanced Training Techniques in Sequential Recommendations: A Comparative Study on Efficiency and Performance

### v1

@@ 1.1.6 Universal Gradient Method for Stochastic Optimization
For any $t\in\mathbb{R}$, by $[t]_{+}:=\max\{t,0\}$, its positive part is denoted. For random variables $X$ and $\xi$, by $\mathbb{E}_{\xi}[X]$ and $\mathbb{E}[X]$, the expectation of $X$ w.r.t. $\xi$, and the full expectation of $X$, respectively are denoted.
@@ 2 Odds Ratio Preference Optimization
Odds Ratio Preference Optimization (ORPO), Hong et al. (2024) incorporates an odds ratio-based penalty to the conventional negative log-likelihood (NLL) loss for differentiating the generation styles between favored and disfavored responses.
F an input sequence $x$, the average log-likelihood of generating the output sequence $y$, of length $m$ tokens, is computed as in the Equation 1. The odds of generating the output sequence $y$ given an input sequence $x$ are defined in the Equation 2:
$\log P_{\theta}(y|x)=\frac{1}{m}\sum_{t=1}^{m}\log P_{\theta}(y_{t}|x,y_{<t})$ (1)
@@
Intuitively, $\textbf{odds}_{\theta}(y|x)=k$ implies that it is $k$ times more likely for the model $\theta$ to generate the output sequence $y$ than not generating it. Thus, the odds ratio of the chosen response $y_{w}$ over the rejected response $y_{l}$, $\textbf{OR}_{\theta}(y_{w},y_{l})$, indicates how much more likely it is for the model $\theta$ to generate $y_{w}$ than $y_{l}$ given input $x$, defined in Equation 3.
@@ 2.1 Objective Function of ORPO
The objective function of ORPO in Equation 4 consists of two components: 1) supervised fine-tuning (SFT) loss ($\mathcal{L}_{SFT}$); 2) relative ratio loss ($\mathcal{L}_{OR}$).
@@
$\mathcal{L}_{SFT}$ follows the conventional causal language modeling negative log-likelihood (NLL) loss function to maximize the likelihood of generating the reference tokens. $\mathcal{L}_{OR}$ in Equation 5 maximizes the odds ratio between the likelihood of generating the disfavored response $y_{w}$ and the disfavored response $y_{l}$. The log odds ratio are wrapped with the log sigmoid function so that $\mathcal{L}_{OR}$ could be minimized by increasing the log odds ratio between $y_{w}$ and $y_{l}$.
@@ 2.2 Gradient of ORPO
The gradient of $\mathcal{L}_{Ratio}$ further justifies using the odds ratio loss. It comprises two terms: one that penalizes the wrong predictions and one that contrasts between chosen and rejected responses, denoted in Equation 6for $d=(x,y_{l},y_{w})\sim D$.
@@
When the odds of the favored responses are relatively higher than the disfavored responses, $\delta(d)$ in Equation 7 will converge to 0. This indicates that the $\delta(d)$ will play the role of a penalty term, accelerating the parameter updates if the model is more likely to generate the rejected responses.
Meanwhile, $h(d)$ in Equation 8 implies a weighted contrast of the two gradients from the chosen and rejected responses. Specifically, $1-P(y|x)$ in the denominators amplifies the gradients when the corresponding side of the likelihood $P(y|x)$ is low. For the chosen responses, this accelerates the model’s adaptation toward the distribution of chosen responses as the likelihood increases.
@@ 3.3 Hydra layer
To apply the sequence-to-sequence potential of the Hydra block, we provide a custom Hydra layer that combines the Hydra block with a standard feed-forward network. The main part of our standard architecture consists of Hydra layers. We have found that Hydra layers work more effectively than Mamba layers. Then we use the same PFFN that was introduced in Mamba4Rec Liu et al. (2024)
@@ 3.4.2 Interpretation of results
It is evident that Hydra shows metrics that are similar to those of Mamba, but its latency is four to five times lower. Simultaneously, models such as LlamaRec exhibit superior performance; however, their substantial parameter count may constrain their applicability in real-world situations. Metrics-wise and latency-wise, SSM-based models outperform the baseline SASRec model on average.
@@ 3.5 Application of ORPO to recommender systems
To improve quality and incorporate human preference into the LLM, we used the ORPO technique described above.
@@ 3.5.2 Data
We used the standard LlamaRec format. To construct the text input, we prepend an instruction to describe the task, followed by both history and candidate items, represented by their titles. Our prompt template is: ” Instruction: Given the user’s history in chronological order, recommend an item from the candidate pool with its index letter; Input: User history: history ; Candidate pool: candidates ; Answer: label ” where history, candidates and label are replaced by history item titles, candidate item titles and the labeld of each data item. For inference, the label position is left empty for the model to provide predictions.
@@ 3.5.3 Experimental setup
For the ORPO procedure we used lr = 5e-6, alpha = 0.05, and an inverse sqrt scheduler with 100 warmup steps, gradient accumulation step equal to 1 and 5 training epochs. We considered 2 datasets - Movielens-1M and Amazon VideoGames.
@@ 3.6.1 Algorithm
We were able to apply the concept of adaptive batching to the aforementioned algorithm (USGM) and enhance its performance in training SSMs in the sequential recommendations domain.
@@
Since the number of iterations of the algorithm with adaptive batching and of those, with which it was compared, differs, since the equality $N_{\text{iterations}}=\frac{\text{Dataset sise}}{\text{Batch size}}=\textit{const}$ is not true for the proposed algorithm, a scaled measure, which is proportional to the epoch number is used as the X axis in the comparison plot, displayed in the Figure 3.
The experiments conducted using the 2Mamba4Rec pipeline, introduced in this paper, on the MovieLens-100k dataset have shown promising results regarding the innovative adaptive batching algorithm, applied to the USGM optimization algorithm by Rodomanov et al. (2024), when compared with the well-known Adam optimizer by Kingma & Ba (2017).
@@ 4 Discussion
Limitations. ORPO and LLM related procedures require a lot of computational resources, especially for larger models. LLM inference itself also requires optimization. SSM based models have poor CPU performance due to their structure and implementation. Furthermore, optimizer that we proposed has poor robustness among its hyperparameters.

### v2

@@ 1 Introduction
Modern digital products are heavily dependent on the performance of the recommender systems, that are frequently utilized to predict future interactions of customers based on their historical behaviour.
While recurrent neural networks (RNNs) and convolutional neural networks (CNNs) led the way in the use of neural networks in sequential recommendation, they suffer from the forgetting issue Kirkpatrick et al. (2017). Recently, researches were able to introduce the Transformer-based models into the field Li et al. (2023).
Due to the quadratic computational cost, these attention-based approaches typically suffer from the inference inefficiency problem despite their impressive performance.
SSM models have been adopted as a replacement for RNNs and CNNs, with Liu et al. (2024) being an example. We develop this idea further, while introducing a novel state-of-the-art sequential recommender model Hydra4Rec, based on the architecture proposed in Hwang et al. (2024).
@@ 1.1.5 Mamba applications to sequential recommendations
1.2 Hydra
Hydra fully utilize the matrix mixer framework, to explore a novel bidirectional sequence mixer. To achieve linear complexity, special class of quasiseparable matrices is considered. A matrix M is quasiseparable if every label element $m_{ij}$ satisfies:
$m_{ij}=\begin{cases}\overrightarrow{\textbf{c}^{T}_{i}}\overrightarrow{\textbf{A}^{\times}_{i:j}}\overrightarrow{\textbf{b}_{j}},&\text{if }i>j\\ \delta_{i},&\text{if }i=j\\ \overleftarrow{\textbf{c}^{T}_{j}}\overleftarrow{\textbf{A}^{\times}_{j:i}}\overleftarrow{\textbf{b}_{i}},&\text{if }i<j\\ \end{cases},$ (1)
where each $\delta_{i}$ is a scalar, $\mathbf{b}_{i},\mathbf{c}_{i}\in\mathbb{R}^{N\times 1}$, and $\mathbf{A}_{i}\in\mathbb{R}^{N\times N}$. Quasiseparable matrices can be used as matrix mixers, this framework is named as Hydra, achieving $O(L)$ complexity. Based on an important property of quasiseparable matrices - quasiseparable matrices can be expressed as a combination of two semiseparable matrices - a subquadratic matrix multiplication algorithm was developed. Then, as a semiseparable matrix structure includes SSMs Dao & Gu (2024), any SSM could be combined with the Hydra framework, resulting in a new efficient bidirectional SSM architecture.
@@ 1.2.1 Universal Gradient Method for Stochastic Optimization
For any $t\in\mathbb{R}$, by $t]_{+}:=\max\{t,0\}$, its positive part is denoted. For random variables $X$ and $\xi$, by $\mathbb{E}_{\xi}[X]$ and $\mathbb{E}[X]$, the expectation of $X$ w.r.t. $\xi$, and the full expectation of $X$, respectively are denoted.
@@ 2 Odds Ratio Preference Optimization
Odds Ratio Preference Optimization (ORPO), Hong et al. (2024) - is novel preference optimization framework that consolidates an odds ratio-based penalty.
The odds of generating the output sequence $y$ given an input sequence $x$ are defined by:
$\log P_{\theta}(y|x)=\frac{1}{m}\sum_{t=1}^{m}\log P_{\theta}(y_{t}|x,y<t)$ (2)
@@
The odds ratio of the chosen response $y_{w}$ over the rejected response $y_{l}$, $\textbf{OR}_{\theta}(y_{w},y_{l})$, indicates how much more likely it is for the model $\theta$ to generate $y_{w}$ than $y_{l}$ given input $x$.
@@ 2.1 Objective Function of ORPO
The objective function of ORPO consists of two components: 1) supervised fine-tuning (SFT) loss ($\mathcal{L}_{SFT}$); 2) relative ratio loss ($\mathcal{L}_{OR}$).
@@
$\mathcal{L}_{SFT}$ follows the conventional causal language modelling negative log-likelihood (NLL) loss function to maximise the probability of generating the reference tokens. The $\mathcal{L}_{OR}$ in the equation maximises the odds ratio between the likelihood of generating the unfavourable response $y_{w}$ and the favourable response $y_{l}$.
@@ 2.2 Gradient of ORPO
The gradient of $\mathcal{L}_{Ratio}$ further justifies the use of the odds ratio loss. It consists of two terms: one that penalises wrong predictions and one that contrasts between selected and rejected answers, for $d=(x,y_{l},y_{w})\sim D$.
@@
If the probabilities of the favoured responses are relatively higher than the disfavoured responses, $\delta(d)$ in the equation will converge to 0. This indicates that the $\delta(d)$ will play the role of a penalty term, speeding up the parameter updates when the model is more likely to generate the disfavoured responses.
Meanwhile, $h(d)$ implies a weighted contrast of the two gradients from the chosen and rejected responses. Specifically, 1P(y|x)1-P(y|x) in the denominators amplifies the gradients when the corresponding side of the likelihood $P(y|x)$ is low. As the likelihood increases, the model adapts more quickly to the distribution of selected responses.
@@ 3.3 Hydra layer
To apply the sequence-to-sequence potential of the Hydra block Hwang et al. (2024), we provide a custom Hydra layer that combines the Hydra block with a standard feed-forward network. The main part of our standard architecture consists of Hydra layers. We have found that Hydra layers work more effectively than standard Mamba layers in several cases. Then we use the same PFFN that was introduced in Mamba4Rec Liu et al. (2024)
@@ 3.4 Benchmarks
All experiments were conducted using the same initial hyperparameters for better interpretability.
@@ 3.4.2 Interpretation of results
It is evident that Hydra shows metrics that are similar, sometimes better to those of Mamba, but its latency is four to five times lower. Simultaneously, models such as LlamaRec exhibit superior performance; however, their substantial parameter count may constrain their applicability in real-world situations. Metrics-wise and latency-wise, SSM-based models outperform the baseline SASRec model on average.
@@ 3.5 Application of ORPO to recommender systems
To improve quality and incorporate human preference into the LLM, we used the ORPO technique described above. Main motivation is ability to slightly improve quality of the large model, showing the potential of preference optimization techniques in recommeder systems.
@@ 3.5.2 Data
We used the standard LlamaRec format. To construct the text input, we prepend an instruction to describe the task, followed by both history and candidate items, represented by their titles. Our prompt template is: ” Instruction: Given the user’s history in chronological order, recommend an item from the candidate pool with its index letter; Input: User history: history ; Candidate pool: candidates ; Answer: label ” where history, candidates and label are replaced by history item titles, candidate item titles and the labeled of each data item. For inference, the label position is left empty for the model to provide predictions.
@@ 3.5.3 Experimental setup
For the ORPO procedure we used lr = 8e-6, ORPO uses very low learning rates compared to traditional SFT or even DPO. This value of 8e-6 comes from the original paper. beta = 0.1 An appendix from the original paper shows how it’s been selected with an ablation study. We use paged adamw optimizer with gradient accumulation steps = 4. We train for 1 training epoch and evaluate each 0.2 steps, also we set warmup steps to 10. We use linear scheduler type. Other parameters are dependant on capabilities of your resources. Training was proceed on 1 A100 80GB videocard.
@@ 3.6.1 Algorithm
We were able to apply the concept of adaptive batching to the aforementioned algorithm (USGM) and enhance its performance in training SSMs in the sequential recommendations domain. The approach presented in this section is novel, since adaptive batching techniques have never been applied to the Rodomanov et al. (2024).
@@ 4 Discussion
Limitations. ORPO and LLM related procedures require a lot of computational resources, especially for larger models. LLM inference itself also requires optimization, but this does not make them unusable in real tasks. SSM based models might have poor CPU performance due to their structure and implementation. Furthermore, optimizer that we proposed has poor robustness among its hyperparameters.